OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick a yellow marble?

OpenStudy (thatgingeronfire):

4 + 3 + 3= 10 total marbles. Probability of picking a red marble is 4/10 or 2/5. Then picking a yellow one would be 3/9 or 1/3. So do \[2/5 \times 1/3= 2/15\].

OpenStudy (thatgingeronfire):

Does that make sense?

OpenStudy (anonymous):

yes can you help me with these one? A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a black marble, replacing it, and then picking a blue marble?

OpenStudy (thatgingeronfire):

Okay, so we have to find the total marbles. 5 + 6 + 4= 15. So to pick a black marble is 4/15. If you replace it the total is still 15, and getting a blue marble would be 6/15. So we do \[4/15\times6/15=8/75\]

OpenStudy (anonymous):

thanks for your help im starting to get it

OpenStudy (thatgingeronfire):

Good! Glad I could help.